OpenStudy (anonymous):

How to solve cos( 11 pi/4

OpenStudy (jdoe0001):

you're asked for the angle?

OpenStudy (jdoe0001):

I mean, there's nothing to solve per se

OpenStudy (anonymous):

It's asking for the exact value of the trigonometric function

OpenStudy (anonymous):

convert 11pi/4 rads into degrees. 11*180/4 = 495 deg 495 deg = 135 deg 135 deg is in Quadrant II on the unit circle 135 deg will create a triangle of 45 deg when going down to the x-axis Now you have a 45-45-90 triangle whose sides must be sqrt(2)/2 for the adj. sqrt(2)/2 for the opp. 1 for the hypotenuse. Cosine = adj/hyp Cos = sqrt(2)/2 But since it's in the second quadrant its adj. side is negative so cos(11pi/4) = -sqrt(2)/2

OpenStudy (anonymous):

That's exactly what I needed, thank you. I didn't realize it created a 45-45-90- triangle. Thank you.